# Training Foundation Models on a Full-Stack AMD Platform: Compute, Networking, and System Design

### v1

@@ V-C1 Context Parallelism
In the backward pass, for the point to point exchange of chunk-boundary tokens, we need only replace every send operation with a receive operation and vice versa. Then we can simultaneously compute the vector Jacobian products for the two convolution operations on each rank. For the value projection, we replace the all-gather with a reduce-scatter in the backward pass and differentiate through the projection. In both forward and backward passes we ensure that the communication overhead remains minimal.
Figure 11: The context parallelism design used to train ZAYA1-base on longer context lengths.
@@ V-D Checkpointing
Our training employs Muon for specific parameter subsets and AdamW for the remainder. Muon optimizes the CCA convolutions (both 1D convs), tied embeddings, residual-scaling vectors, all layer norms (RMSNorms, excluding qk-norm), and CCA temperatures. All remaining parameters—Q/K/V projections, output projections, router MLPs, down-projections, and MoE experts—are optimized with AdamW. We model the distributed checkpointing procedure as follows: Let $P$ be the total number of trainable parameters saved, with $P=P_{M}+P_{A}$, where $P_{M}$ and $P_{A}$ are the numbers of parameters optimized with Muon and AdamW, respectively. Let $b_{\mathrm{lp}}$ be the bytes per low-precision weight, $b_{\mathrm{hp}}$ the bytes per high-precision value ($b_{\mathrm{lp}}=2$ for bf16 and $b_{\mathrm{hp}}=4$ for fp32), dp_degree the ZeRO-1 data-parallel degree, and $m_{r}$ a small per-rank metadata memory overhead in bytes (RNG states, schedulers, JSON, etc.).
@@ Appendix D Communication Results
While there wasn’t sufficient room in the main paper, we include the AlltoAll communication operation results in this appendix. Both inter-node and intra-node are included. The expert block of ZAYA1 was able to fit within HBM, but subsequent models will be too large, and expert parallelism will need to be used. Therefore, we intend to co-design the expert parallelism implementation of our training framework with the underlying intra-node fabric in order to alleviate the costly AlltoAll operation when shuffling tokens to/from experts in every layer.

### v2

@@ V-C1 Context-Parallelism
In the backward pass, for the point to point exchange of chunk-boundary tokens, we need only replace every send operation with a receive operation and vice versa. Then we can simultaneously compute the vector Jacobian products for the two convolution operations on each rank. For the value projection, we replace the AllGather with a ReduceScatter in the backward pass and differentiate through the projection. In both forward and backward passes we ensure that the communication overhead remains minimal.
Figure 11: The context-parallelism design used to train ZAYA1-base on longer context lengths.
@@ V-D Checkpointing
Our training employs Muon for specific parameter subsets and AdamW for the remainder. AdamW optimizes the CCA convolutions (both 1D convs), tied embeddings, residual-scaling vectors, all layer norms (RMSNorms, excluding qk-norm), and CCA temperatures. All remaining parameters—Q/K/V projections, output projections, router MLPs, down-projections, and MoE experts—are optimized with Muon. We model the distributed checkpointing procedure as follows: Let $P$ be the total number of trainable parameters saved, with $P=P_{M}+P_{A}$, where $P_{M}$ and $P_{A}$ are the numbers of parameters optimized with Muon and AdamW, respectively. Let $b_{\mathrm{lp}}$ be the bytes per low-precision weight, $b_{\mathrm{hp}}$ the bytes per high-precision value ($b_{\mathrm{lp}}=2$ for bf16 and $b_{\mathrm{hp}}=4$ for fp32), dp_degree the ZeRO-1 data-parallel degree, and $m_{r}$ a small per-rank metadata memory overhead in bytes (RNG states, schedulers, JSON, etc.).
@@ Appendix D Communication Results
While there wasn’t sufficient room in the main paper, we include the All-to-All communication operation results in this appendix. Both inter-node and intra-node are included. The expert block of ZAYA1 was able to fit within HBM, but subsequent models will be too large, and expert-parallelism will need to be used. Therefore, we intend to co-design the expert-parallelism implementation of our training framework with the underlying intra-node fabric in order to alleviate the costly All-to-All operation when shuffling tokens to/from experts in every layer.